OpenStudy (gabylovesyou):

What are the possible rational zeros of f(x) = x4 + 6x3 − 3x2 + 17x − 15? ± 1, ± 3, ± 5, ± 15 ± 1, ± 1 over 3, ± 1 over 5, ± 1 over 15 ± 1, ± 3, ± 6, ± 15, ± 17 ± 1, ± 1 over 3, ± 1 over 6, ± 1 over 15, ± 1 over 17

OpenStudy (gabylovesyou):

@Nnesha

OpenStudy (fibonaccichick666):

so did you graph it?

Nnesha (nnesha):

`Rational Roots Test` we have to find possible zeros by testing each number from p/q fraction \[\rm \frac{ p }{ q }=\frac{ \pm factors~of~constant~term }{\pm factors~of~leading ~coefficient }\]

Nnesha (nnesha):

so what's the leading coefficient and constant term in that equation ? or as mentioned above you can just graph it zeros (x-intercept ,solution) are points where line intersect the x-axis

OpenStudy (gabylovesyou):

um... i have no idea ;/ I'm so lost in this.

OpenStudy (gabylovesyou):

how do i find the factors

Nnesha (nnesha):

well i just noticed the graph will give you EXACT solution but we need POSSIBLE

Nnesha (nnesha):

so go with p/q method what is the leading coefficient in that equation ? leading coefficient = coefficient of highest degree variable

Nnesha (nnesha):

and what is the constant term ?

OpenStudy (gabylovesyou):

x^4

OpenStudy (gabylovesyou):

has the highest degree..

Nnesha (nnesha):

that's same as 1x^4 so one is leading coefficient

OpenStudy (gabylovesyou):

ok

Nnesha (nnesha):

and constant term ?

OpenStudy (gabylovesyou):

what is that ?

Nnesha (nnesha):

constant term ( number without any variable )

OpenStudy (gabylovesyou):

15

OpenStudy (gabylovesyou):

or would it be -15 ?

OpenStudy (gabylovesyou):

@Nnesha

Nnesha (nnesha):

sorry my computer was shutdown and yes that's correct

Nnesha (nnesha):

what are the factors of 15 ?

OpenStudy (gabylovesyou):

1,3,5,15

Nnesha (nnesha):

right \[\rm \frac{ p }{ q }=\frac{ \pm1, \pm 3, \pm 3 , \pm5 , \pm 15 }{\pm 1 }\] now remember each number at the numerator is dividing by 1 simplify each fraction to eliminate the identical values

Nnesha (nnesha):

that's it!

OpenStudy (gabylovesyou):

so its A ?

Nnesha (nnesha):

that's correct

OpenStudy (gabylovesyou):

Thank you

Nnesha (nnesha):

yw